OpenStudy (anonymous):

How would I prove 2/2^n = 1/2^n-1

OpenStudy (anonymous):

\[\begin{align*}\frac{2}{2^n}&=\frac{1}{2^{n-1}}\\ \frac{2^1}{2^n}&=\cdots\\ 2^{1-n}&=\cdots\\ \frac{1}{2^{-(1-n)}}&=\cdots\\ \frac{1}{2^{n-1}}&=\cdots \end{align*}\]

OpenStudy (dumbcow):

or in other words...you reduce the fraction by a factor of 2 just like 2/4 = 1/2

OpenStudy (dan815):